# CyberSOCEval: Benchmarking LLMs Capabilities for Malware Analysis and Threat Intelligence Reasoning

### v1

@@ 1 Introduction
The growing complexity and volume of cyber threats have overwhelmed cyber defenders working in Security Operations Centers (SOCs). Large Language Models (LLMs) offer a promising avenue for automation, yet currently have skills gaps that hinder their utility, as we show below. Evaluations like those introduced here can guide model developers in improving their models’ cyber capabilities to better help cyber defenders. Simultaneously, threat actors are using Artificial Intelligence (AI) to enhance their cyber attacks. LLMs can generate highly convincing phishing emails, messages, and fake websites, expanding beyond attackers’ native languages. AI systems also accelerate coding, serve as valuable technical assistants, and can aid in the research of evasion techniques. Malware developers benefit from these systems just as legitimate engineers do. Attackers have been observed to leverage all of these new AI capabilities in the wild, and we can expect continued attacker adoption of AI as capabilities grow (Machine Learning-Based Malware Detection in a Production Setting; sec. 5.5.2 Krasser et al. (2025)).
@@ 2 Related Work
Sophos (2024) included three fundamental SOC tasks to benchmark LLMs: (1) incident investigation (converting natural-language questions about telemetry into SIEM queries), (2) incident summarization from raw SOC data, and (3) incident severity rating. The Sophos team did not open source their evals, but their technical report underscores the need for public SOC benchmarks, a gap that CyberSOCEval begins to help close.
The Oniagbi (2024) study of LLM agents uplift for Tier-1 SOC triage in a synthetic lab setting found that LLM assistance significantly accelerated initial triage and reduced analyst cognitive load, yet worked best as a co-pilot rather than an autonomous agent. This study also did not include an open source evaluation framework.
@@
Bhusal et al. (2024) targets cyber defense skills in the Industrial Control Systems domain, and does include an open benchmark; but because it focuses on a single industry vertical, it does not address the need for a more general benchmark suite.
CTIBench by  Alam et al. (2024) is a recent benchmark explicitly designed for cyber threat intelligence (CTI) tasks. It introduces a suite of four complementary evaluations covering threat-intel question answering and analytical reasoning: multiple-choice questions to probe threat knowledge, root cause mapping of vulnerabilities to underlying causes, vulnerability severity prediction using standards like CVSS, and threat actor attribution given attack descriptions.
Ji et al. (2024) is a bilingual (English and Chinese) framework focused on threat intelligence and incident analysis. SEvenLLM curated a high-quality corpus of 8,485 threat reports to fine-tune domain-specific LLMs on 28 cybersecurity tasks, and compiled a 1,300-sample QA benchmark to evaluate CTI analytical capabilities.
@@
Beyond the specialized areas above, there have been broader efforts to benchmark LLMs on security knowledge and defensive reasoning. Wan et al. (2024) took a two-pronged approach: it assessed LLMs’ vulnerabilities (e.g. propensity for insecure code and malicious instruction-following) and their dual use capabilities like vulnerability identification.
Other works have constructed extensive question-answer banks to test factual security knowledge: for example, CyberMetric by Tihanyi et al. (2024) generated thousands of multiple-choice questions from standards (NIST, CWE, etc.) and evaluated dozens of models on this wide-ranging quiz.
@@ 4.1 Malware Analysis
For each model evaluated against the Malware Analysis benchmark, Figure 8 presents the share of questions that the model answered correctly broken out by question difficulty rating. The overall accuracy scores range from approximately 15-28%, which should be compared to a baseline of approximately 0.63% accuracy that would be expected from purely random guessing (see Appendix B).
@@
Given the large size of the Hybrid Analysis reports, which was often near the 128,000 tokens context window limit of several models under test, we also explored whether the entirety of the Hybrid Analysis report was needed to achieve the given performance level. We found that in general, an abridged version of the report does not meaningfully affect performance, and so models with smaller context windows may be evaluated on this benchmark as well, by setting a flag to filter the input report. More detail on the logic used to reduce the report size can be found in Appendix D, along with evidence that this filtering has very little effect on benchmark performance.
@@ 5.2 Multimodal Threat Intelligence Reasoning
The inclusion of multimodal threat intelligence reports (e.g. combining textual IOCs with tables or diagrams) in this benchmark aims to address a gap over existing frameworks like CTIBench by  Alam et al. (2024) and SEvenLLM by Ji et al. (2024). Our approach better reflects the reality of how threat intelligence is communicated in practice, where critical information is often distributed across text, tables, diagrams, and other visual elements.
@@
Figure 11: Image included in threat report at CrowdStrike (2025) which is needed to correctly answer a benchmark question about the initial access vector used by the threat actor. All six of the models under test answer correctly when provided with the report in image form, while only one model (o3) is able to answer correctly when receiving only text input. Interestingly, when models are given both text and image input simulataneously, only half of them (o3, 4o, and gemini-2.5-pro) answer correctly, indicating that the models may be overly reliant on text inputs when provided multiple modalities.

### v2

@@ 1 Introduction
The growing complexity and volume of cyber threats have overwhelmed cyber defenders working in Security Operations Centers (SOCs). Large Language Models (LLMs) offer a promising avenue for automation, yet currently have skills gaps that hinder their utility, as we show below. Evaluations like those introduced here can guide model developers in improving their models’ cyber capabilities to better help cyber defenders. Simultaneously, threat actors are using Artificial Intelligence (AI) to enhance their cyber attacks. LLMs can generate highly convincing phishing emails, messages, and fake websites, expanding beyond attackers’ native languages. AI systems also accelerate coding, serve as valuable technical assistants, and can aid in the research of evasion techniques. Malware developers benefit from these systems just as legitimate engineers do. Attackers have been observed to leverage all of these new AI capabilities in the wild, and we can expect continued attacker adoption of AI as capabilities grow (Machine Learning-Based Malware Detection in a Production Setting; sec. 5.5.2 Krasser2025malware).
@@ 2 Related Work
sophos-benchmark included three fundamental SOC tasks to benchmark LLMs: (1) incident investigation (converting natural-language questions about telemetry into SIEM queries), (2) incident summarization from raw SOC data, and (3) incident severity rating. The Sophos team did not open source their evals, but their technical report underscores the need for public SOC benchmarks, a gap that CyberSOCEval begins to help close.
The oniagbi-evaluation study of LLM agents uplift for Tier-1 SOC triage in a synthetic lab setting found that LLM assistance significantly accelerated initial triage and reduced analyst cognitive load, yet worked best as a co-pilot rather than an autonomous agent. This study also did not include an open source evaluation framework.
@@
bhusal2024securebenchmarkinglargelanguage targets cyber defense skills in the Industrial Control Systems domain, and does include an open benchmark; but because it focuses on a single industry vertical, it does not address the need for a more general benchmark suite.
CTIBench by  alam2024ctibenchbenchmarkevaluatingllms is a recent benchmark explicitly designed for cyber threat intelligence (CTI) tasks. It introduces a suite of four complementary evaluations covering threat-intel question answering and analytical reasoning: multiple-choice questions to probe threat knowledge, root cause mapping of vulnerabilities to underlying causes, vulnerability severity prediction using standards like CVSS, and threat actor attribution given attack descriptions.
ji2024sevenllmbenchmarkingelicitingenhancing is a bilingual (English and Chinese) framework focused on threat intelligence and incident analysis. SEvenLLM curated a high-quality corpus of 8,485 threat reports to fine-tune domain-specific LLMs on 28 cybersecurity tasks, and compiled a 1,300-sample QA benchmark to evaluate CTI analytical capabilities.
@@
Beyond the specialized areas above, there have been broader efforts to benchmark LLMs on security knowledge and defensive reasoning. wan2024cyberseceval3advancingevaluation took a two-pronged approach: it assessed LLMs’ vulnerabilities (e.g. propensity for insecure code and malicious instruction-following) and their dual use capabilities like vulnerability identification.
Other works have constructed extensive question-answer banks to test factual security knowledge: for example, CyberMetric by cybermetric generated thousands of multiple-choice questions from standards (NIST, CWE, etc.) and evaluated dozens of models on this wide-ranging quiz.
@@ 4.1 Malware Analysis
For each model evaluated against the Malware Analysis benchmark, Figure 8 presents the share of questions that the model answered correctly broken out by question difficulty rating. The overall accuracy scores range from approximately 23-34%, which should be compared to a baseline of approximately 0.63% accuracy that would be expected from purely random guessing (see Appendix B).
@@
Given the large size of the Hybrid Analysis reports, which was often near or in excess of the 128,000 tokens context window limit of several models under test, we also explored whether the entirety of the Hybrid Analysis report was needed to achieve the given performance level. We found that in general, an abridged version of the report does not meaningfully affect performance, and so models with smaller context windows may be evaluated on this benchmark as well, by setting a flag to filter the input report. More detail on the logic used to reduce the report size can be found in Appendix D, along with evidence that this filtering has very little effect on benchmark performance.
@@ 5.2 Multimodal Threat Intelligence Reasoning
The inclusion of multimodal threat intelligence reports (e.g. combining textual IOCs with tables or diagrams) in this benchmark aims to address a gap over existing frameworks like CTIBench by  alam2024ctibenchbenchmarkevaluatingllms and SEvenLLM by ji2024sevenllmbenchmarkingelicitingenhancing. Our approach better reflects the reality of how threat intelligence is communicated in practice, where critical information is often distributed across text, tables, diagrams, and other visual elements.
@@
Figure 11: Image included in threat report at csit-2311 which is needed to correctly answer a benchmark question about the initial access vector used by the threat actor. All six of the models under test answer correctly when provided with the report in image form, while only one model (o3) is able to answer correctly when receiving only text input. Interestingly, when models are given both text and image input simulataneously, only half of them (o3, 4o, and gemini-2.5-pro) answer correctly, indicating that the models may be overly reliant on text inputs when provided multiple modalities.
@@ 7 Acknowledgements
Additionally, the authors are extremely grateful to Anthony Penta, Ram Shankar Siva Kumar, and the Microsoft Security AI cyber evaluation group for reporting a bug in the Malware Analysis benchmark code. Based on their findings, we have corrected the bug, and updated the results in the paper for this benchmark in October 2025.